If logk4=2/3 then k=?
Thank you

Answer by stanbon(75887)

If logk4=2/3 then k=?
K^(2/3)=4
Raise both sides to the 3/2 power to get:
k=4^(3/2)
k=2^3
k=8
Cheers,
Stan H.